DIFFUSION MONTE CARLO CALCULATIONS OF MINIMUM ENERGY PATHS FOR THE ISOTOPIC VARIANTS OF THE CH_{3}^{+} + H_{2} ↔ CH_{5}^{+} ↔ CH_{3}^{+} + H_{2} REACTION

\underline{CHARLOTTE E. HINKLE}, ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Protonated methane is of interest to astrochemists due to its presumed importance as a reaction intermediate in the reaction involving CH_{3}^{+} + HD within the interstellar medium. However, within the interstellar medium there is a nonstatistical H/D isotopic abundance observed for the isotopologues of CH_{3}^{+}. Kinetic studies performed by Gerlich and co-workers determined that the reactions

CH_{3−n}D_{n}^{+} + HD → CH_{4−n}D_{n+1}^{+} → CH_{2−n}D_{n+1}^{+} + H_{2}  (1)

have identical net rate constants regardless of the value of n.^{a} We have calculated zero-point corrected energies and wave functions of the CH_{3}^{+} + H_{2} system^{b} and its deuterated analogs as functions of the center of mass separation between CH_{3}^{+} and H_{2}, R. We can divide these simulations into distinct ranges of R; long-range interactions, complexation, and intermediate distances. Analysis of the wave functions associated with these three ranges of R allows us to study how zero-point energy influences the approach geometries that are sampled during low-energy collisions.

aO. Asvany, S. Schlemmer, D. Gerlich, Astrophys. J. 617, 685 (2004).

bC. E. Hinkle, A. B. McCoy, J. Phys. Chem. Lett. 1, 562 (2010).